simultaneous equation
7x-2y=26 (1)
6x+5y=29

Found 2 solutions by mananth, Gogonati:
Answer by mananth(16946)

7x-2y=26 .............1
6x+5y=29 .............2

multiply (1)by 5
Multiply (2) by 2
35x-10y =130
12x+10y =58

Add the two equations

47x=188

/47

x=4

plug value of x in (1)

7x-2y=26
28-2y=26
-2y=26-28
-2y=-2
y=1

Answer by Gogonati(855)

y=7x-26/2
6x+5(7x-26)/2=29
x=159/47
y=109/94